OpenStudy (anonymous):

The volume of a gas is 200.0 mL at 350.0 kPa pressure. What will the volume be when the pressure is reduced to 120.0 kPa, assuming the temperature remains constant?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

@zakrocks007 but how did you get that, i dont think thats right

OpenStudy (anonymous):

PV always remain constant

OpenStudy (anonymous):

so when P is reduced, to compensate you must increase volume

OpenStudy (anonymous):

350 * 200=a 120* x=a

OpenStudy (anonymous):

350 * 200=120* x you can solve for x